# Intelligent System for Automated Molecular Patent Infringement Assessment

### v1

@@ Abstract
Automated drug discovery offers significant potential for accelerating the development of novel therapeutics by substituting labor-intensive human workflows with machine-driven processes. However, a critical bottleneck persists in the inability of current automated frameworks to assess whether newly designed molecules infringe upon existing patents, posing significant legal and financial risks. We introduce PatentFinder, a novel tool-enhanced and multi-agent framework that accurately and comprehensively evaluates small molecules for patent infringement. It incorporates both heuristic and model-based tools tailored for decomposed subtasks, featuring: MarkushParser, which is capable of optical chemical structure recognition of molecular and Markush structures, and MarkushMatcher, which enhances large language models’ ability to extract substituent groups from molecules accurately. On our benchmark dataset MolPatent-240, PatentFinder outperforms baseline approaches that rely solely on large language models, demonstrating a 13.8% increase in F1-score and a 12% rise in accuracy. Experimental results demonstrate that PatentFinder mitigates label bias to produce balanced predictions and autonomously generates detailed, interpretable patent infringement reports. This work not only addresses a pivotal challenge in automated drug discovery but also demonstrates the potential of decomposing complex scientific tasks into manageable subtasks for specialized, tool-augmented agents.
@@ 1 Introduction
Drug discovery [1, 2] remains a cornerstone in humanity’s ongoing fight against disease, representing a vital pathway to improving global health outcomes. Traditional drug discovery [3] methods rely on wet-lab experiments and human expertise, rendering the process labor-intensive and time-consuming. To enhance efficiency and expedite the development of novel therapeutics, prior studies envision an automated drug discovery [4] pipeline, replacing manual workflows with machine-driven processes. This transition has shown remarkable potential in small-molecule drug discovery. Specifically, the advances in retrosynthesis planning [5, 6, 7, 8, 9], wet lab experiment automation [10, 11, 12, 13, 14], molecular docking [15, 16, 17, 18, 19], and de novo drug design [20, 21, 22, 23, 24] have demonstrated the feasibility of reducing human intervention while enhancing throughput and precision.
Despite these developments, AI models that learn existing drug data and structures to create new molecules may use protected compound structures during generation, leading to substantial legal and financial risks [25]. Such risk exists even for novel molecules, as they may still fall within the scope of existing patents due to overlapping Markush definitions or structural analogies. Current patentability assessments [26, 27] still heavily rely on manual inspection by pharmaceutical chemistry experts, which is labor-intensive, costly, and difficult to scale. This limitation hinders the realization of fully integrated, closed-loop automation, underscoring the need for an intelligent, end-to-end solution to streamline and enhance the patent assessment workflow.
@@
Large language models (LLMs) [28, 29, 30] have demonstrated exceptional capabilities in analysis, reasoning, and problem-solving across diverse tasks [30, 31, 32, 33]. Multimodal LLMs, in particular, extend these capabilities to interpret images [34, 35, 36, 37] and PDF documents [38], making their potential applications in scientific domains, including drug discovery, increasingly practical. However, directly using LLMs for patent infringement assessment reveals several significant obstacles: (1) Inferior reasoning and lack of explainability. Without explicit stepwise guidance, LLMs often generate hallucinations or biased judgments, with reasoning pathways that remain opaque. This lack of reliability and interpretability diminishes their suitability for high-stakes applications such as patent infringement assessment. (2) Limited capability in recognizing molecular structures. As shown in Fig.5, LLMs’ built-in image and PDF parsers struggle to interpret molecular and Markush diagrams in patent claims accurately. This deficiency hampers their ability to analyze essential structural details. (3) Inadequate recognition of substituent groups. Results in Fig.3 demonstrate that, when comparing candidate molecules to Markush definitions, LLMs frequently misinterpret molecular structures, resulting in incorrect predictions for substituent groups. (4) Absence of public patent infringement assessment benchmark. Although there are some open-source patent datasets available [39, 40], current researches lack a public benchmark for patent infringement assessment.
To overcome these challenges, we present PatentFinder, a novel end-to-end framework designed to accurately and comprehensively assess whether a molecule falls within the scope of a given patent. As illustrated in Fig.1, it adopts a tool-enhanced [41, 11, 42] and multi-agent [43, 44, 45] framework that decomposes the complex patent assessment task into manageable subtasks handled by specialized agents. The major innovations of PatentFinder include: (1) The multi-agent framework enables PatentFinder to generate intermediate outputs and compile them into a comprehensive infringement report, thereby enhancing the interpretability of the reasoning process. (2) We design a reverse data construction strategy and trained MarkushMatcher, which is capable of assisting LLMs in accurate substituent recognition; (3) MarkushParser, a Markush structure parsing tool model, which is trained to enable robust interpretation of molecular and Markush structures, even from complex visual representations; (4) To facilitate the evaluation of molecular patent infringement algorithms, we develop MolPatent-240, a molecular patent infringement evaluation benchmark that comprises 240 molecule-patent pairs.
@@
Overall, extensive experiments in this paper demonstrate the following observations: (1) PatentFinder achieves state-of-the-art performance, significantly outperforming baseline approaches that rely solely on LLMs. Specifically, PatentFinder surpassed baselines by 13.8% in F1-score and 12% in accuracy, demonstrating its effectiveness in handling the nuanced demands of this task. (2) As illustrated in Fig.2e, PatentFinder autonomously generates comprehensive patent infringement reports. These reports integrate detailed molecular structure analyses and step-by-step reasoning, resulting in outputs that are both highly informative and easily interpretable. This capability represents a significant step forward in enhancing the transparency and usability of automated patent analysis solutions. Moreover, our approach of decomposing complex tasks into modular components, each managed by a specialized tool agent, holds potential for application in other scientific domains.
@@ 2 Results and Discussion
PatentFinder introduces two critical designs to facilitate the assessment of small-molecule patent infringement: (1) a multi-agent framework based on LLMs, where various agents equipped with tailored tools address relatively simple subtasks, and (2) two neural network models fine-tuned for specific tasks within the patent analysis process, addressing the performance deficiencies of LLMs in these tasks. To evaluate the effectiveness of these designs, we conduct separate experiments for each. For the assessment of PatentFinder’s multi-agent framework, we manually collect 70 real-world patents to construct the MolPatent-240 patent infringement benchmark and compare PatentFinder with state-of-the-art LLMs on small-molecule patent protection discrimination. For the fine-tuned tool models, we evaluate them on the tasks of Markush molecular structure matching and Markush image recognition. The experimental results demonstrate that these tool models effectively compensate for the deficiencies of general-purpose LLMs in Markush understanding tasks.
@@ 2.1.1 Experiment Settings
To demonstrate the effectiveness of PatentFinder for patent infringement assessment, we construct a patent protection benchmark – MolPatent-240. To illustrate the rationale behind our agent-based design, we compare PatentFinder with the direct prompting of leading LLMs, which are tasked with reasoning and determining whether a given molecule infringes on a patent without any additional guidance. The baselines include state-of-the-art large language models Gemini-1.5-Pro [38], Claude-3.5-Sonnet 111https://www-cdn.anthropic.com/de8ba9b01c9ab7cbabf5c33b80b7bbc618857627/Model_Card_Claude_3.pdf., GPT-4o [34], and OpenAI-o1 222https://openai.com/index/introducing-openai-o1-preview. We leverage the Chain-of-Thought prompts [46] for all the baseline models to enhance their reasoning ability. As shown in Table 1, we explore three distinct schemes to assess these models for patent infringement assessment: (1) Patent PDF: In this scheme, the patent PDF file is directly provided to the LLM. This approach requires the model to have a built-in PDF parsing capability. (2) Patent Text + Markush Image: The model is fed both the textual content of the patent and the core Markush structure images claimed in the patent. This scheme requires the model to handle both textual and visual information. (3) Patent Text + Markush String: In this case, all Markush images in the patent are first parsed into extended SMILES strings using our MarkushParser model. These Markush strings, along with the patent’s textual content, are then provided to LLMs in text format, allowing us to utilize text-only models in this scheme. Since the Patent PDF and Patent Text + Markush Image schemes require LLMs to handle PDF parsing and image processing, only Gemini, Claude, and GPT-4o are used for these experiments.
@@ 2.1.2 Results & Observations
Experimental results in Table 1 demonstrate that PatentFinder significantly outperforms the direct patent infringement classification with LLMs, regardless of the input scheme adopted. Specifically, our multi-agent collaboration framework achieves a 12.5% improvement in F1-score (relative to Gemini with patent text + Markush string) and an 11% increase in TNR (relative to Gemini with patent text + Markush image) compared to the best-performing baseline models.
@@
Interestingly, we find that some popular LLMs (e.g., GPT-4o and Claude-3.5-Sonnet) tend to adopt an overly conservative classification approach, frequently assuming that the query molecule is protected by the given patent. They exhibit a TPR that is significantly higher than their TNR, often by 50% or more. Notably, GPT-4o (patent text + Markush string) displays a TPR that exceeds its TNR by 69.2%. Given that MolPatent-240 has a roughly balanced positive-to-negative example ratio (1:1.3), this trend suggests that many leading LLMs tend to classify samples as positive indiscriminately, with some models performing worse than random guessing when distinguishing negative samples (TNR).
In contrast, PatentFinder demonstrates robust performance in distinguishing both positive and negative samples. It achieves a high TPR (83.7%) alongside a high TNR (84.6%), showcasing balanced and accurate classification across the MolPatent-240 dataset.
@@ 2.1.3 Case Studies
In real-world molecular patent infringement assessment, an effective and trustworthy collaboration between models and human experts requires the model to produce human-readable infringement analysis reports. These reports should not only deliver a definitive judgment on whether infringement has occurred but also provide a comprehensive and detailed explanation of intermediate reasoning steps. To this end, we evaluate not only the accuracy of the infringement assessment but also the validity of reasoning steps across different methods. Specifically, we compare the intermediate outputs of baseline models and PatentFinder on a real example from the MolPatent-240 dataset. The results are illustrated in Fig.2. For baseline models, we summarize the step-by-step reasoning and final decisions of Gemini, GPT-4o, and Claude, while for PatentFinder, we present the intermediate outputs from each agent and the final infringement report generated by the Planner based on these analyses.
As shown in Fig.2b, directly using LLMs for patent protection determination often leads to hallucinations in intermediate reasoning steps, resulting in incorrect conclusions based on non-existent evidence. For instance, Gemini provides no intermediate analysis and arrives at an incorrect judgment after merely comparing the molecular skeletons to the Markush structure. GPT-4o attempts to match the skeleton against the Markush structure but misidentifies the substituent D1 and introduces a fictitious substituent B. Although Claude reaches the correct conclusion, its analysis erroneously identifies R[21] as a chlorine atom and fails to determine the values of three other substituents.
From the baseline results, two key observations emerge: (1) Only GPT-4o attempts a fine-grained comparison of all substituents, while the other two methods deliver conclusions without detailed analysis. (2) LLMs exhibit limited capabilities in comparing chemical structures, with reasoning errors propagating through subsequent analysis steps.
@@ 2.2 Evaluation of MarkushMatcher
To showcase the ability of MarkushMatcher in delivering precise structural matching results, we curated a dataset comprising 1,000 samples for substituent extraction. The comparison focused on four key metrics: (1) Accuracy: The proportion of predictions that perfectly match the true substituent results. Perfect matches are defined as having a Tanimoto [48, 49] similarity of 100%. (2) Validity: The proportion of valid predictions. A prediction is considered valid if it includes all substituents mentioned in the Markush structure and each predicted substituent corresponds to a legitimate molecular structure. (3) Average Tanimoto Similarity: The average Tanimoto Similarity between each predicted substituent and its corresponding ground truth substituent, computed using Morgan Fingerprints [50] for all substituents in a sample, the higher the better. (4) Levenshtein Distance: The distance between the predicted results and the ground truth is calculated directly at the string level, the lower the better.
@@
We compare MarkushMatcher with structure-matching baselines implemented using the rule-based – RDKit [51] and LLM-based – GPT-4o methods, respectively. As shown in Fig. 3a-c, MarkushMatcher outperforms both baseline methods on the substituent groups prediction task. As a deterministic algorithm, the rule-based approach – RDKit struggles with complex substituent matching tasks. While effective for simple tasks involving single-link substituents, it performs poorly in more intricate scenarios. This limitation resulted in near-zero accuracy and validity scores on our test set (cf. Fig. 3a,b), with unsatisfactory results for the other two metrics as well. Using LLMs alone does not yield satisfactory results either. Although GPT-4o is able to provide chemically valid solutions in many scenarios, which results in the high validity in Fig. 3c, it fails to predict the correct substituent group values in most cases, leading to low similarity, high Leveshtein distance, and limited accuracy. Common errors include: a) Confusing the values of different substituents; b) Incorrectly identifying substituent positions on the Markush structure; and c) Misclassifying scaffold atoms as part of the substituents. MarkushMatcher outperforms the other two methods by a large margin, with an accuracy of 66.8%. In comparison with GPT-4o, MarkushMatcher increases the mean average Tanimoto similarity from 21.3% to 92.9% and has an 8.8 times lower average Levenshtein distance. The training details of MarkushMatcher are presented in Fig. 3d, e. During the model’s training, the training loss converges eventually, with a steady increase in both accuracy and validity. In the training data of MarkushMatcher, monatomic substituents such as C, N, and O are the most frequently occurring, while the heavy atom counts of other groups follow an approximately bell-shaped distribution.
@@
We also find that in certain cases, even with the given Markush and molecular structures, it is impossible to determine a unique result without additional contextual information. For instance, the presence of exchangeable groups (cf. Fig.3f) and adjacent groups (cf. Fig.3g) in the Markush structures can introduce complexities. Specifically, if two substituent groups are connected to the same atom, the correlation between the group identifiers and the substructures may become ambiguous. Similarly, in cases where multiple substituent groups are adjacent, there could be various valid allocation schemes when specifying the group values.

### v2

@@ Abstract
Automated drug discovery offers significant potential for accelerating the development of novel therapeutics by substituting labor-intensive human workflows with machine-driven processes. However, molecules generated by artificial intelligence may unintentionally infringe on existing patents, posing legal and financial risks that impede the full automation of drug discovery pipelines. This paper introduces PatentFinder, a novel multi-agent and tool-enhanced intelligence system that can accurately and comprehensively evaluate small molecules for patent infringement. PatentFinder features five specialized agents that collaboratively analyze patent claims and molecular structures with heuristic and model-based tools, generating interpretable infringement reports. To support systematic evaluation, we curate MolPatent-240, a benchmark dataset tailored for patent infringement assessment algorithms. On this benchmark, PatentFinder outperforms baseline methods that rely solely on large language models or specialized chemical tools, achieving a 13.8% improvement in F1-score and a 12% increase in accuracy. Additionally, PatentFinder autonomously generates detailed and interpretable patent infringement reports, showcasing enhanced accuracy and improved interpretability. The high accuracy and interpretability of PatentFinder make it a valuable and reliable tool for automating patent infringement assessments, offering a practical solution for integrating patent protection analysis into the drug discovery pipeline.
@@ 1 Introduction
Drug discovery [1, 2] remains a cornerstone in humanity’s ongoing fight against disease, representing a vital pathway to improving global health outcomes. Traditional methods [3] rely on wet-lab experiments and human expertise, making the process resource-intensive and time-consuming. Recent advances in artificial intelligence (AI) have been revolutionizing this field with automated drug discovery pipelines. AI-driven solutions have achieved remarkable successes across key tasks, including retrosynthesis planning [4, 5, 6, 7, 8], wet lab experiment automation [9, 10, 11, 12, 13], molecular docking [14, 15, 16, 17, 18], and de novo drug design [19, 20, 21, 22, 23]. These solutions significantly reduce human intervention while enhancing the throughput and precision of drug development workflows. However, the rapid adoption of AI in drug discovery introduces new challenges. While the AI-generated molecules may be novel, they could still fall within the scope of existing patents without intention [24]. This raises substantial legal and financial risks for pharmaceutical companies, underscoring the critical need for an integrated patent infringement assessment within AI-guided drug discovery pipelines. Current patentability assessment workflows [25, 26], which still rely on manual inspections by pharmaceutical chemistry experts, pose a significant bottleneck to achieving fully automated drug discovery. An intelligent, end-to-end system is demanded to streamline and enhance the patent infringement assessment process.
@@
Large language models (LLMs) [27, 28, 29, 30] have demonstrated exceptional capabilities in analysis, reasoning, and problem-solving across diverse tasks [29, 31, 32, 33]. In particular, multimodal LLMs extend these capabilities to interpret images [34, 35, 36, 30] and PDF documents [37], making them increasingly viable for scientific domains like drug discovery. These advancements also highlight the potential for integrating LLMs into patent protection analysis, enabling tasks like extracting critical information from patents and comparing the structures of new molecules against patent claims. However, applying LLMs directly to patent infringement assessment reveals several significant obstacles: (1) Opaque Reasoning Paths: LLMs often yield conservative evaluations that lack reliability and interpretability, making their reasoning pathways difficult to trust in high-stakes applications, like patent infringement assessment. (2) Inaccurate Parsing of Molecular Structures: The built-in image and PDF parsers in LLMs struggle to interpret molecular diagrams and Markush notations in patents accurately. This limitation, likely resulting from insufficient training on molecular data, especially on complex Markush structures, impairs their capability to perform fine-grained molecular analyses that patent claims require. (3) Lack of Benchmarks for Patent Infringement Assessment: Currently, no open-source benchmark is available for patent infringement assessment, although there are some datasets designed for other patent-relevant tasks [38, 39]. The lack of benchmarks limits the systematical evaluation of relevant algorithms.
To address these challenges, we introduce PatentFinder, an innovative end-to-end intelligent system designed to accurately and comprehensively assess whether a molecule falls within the scope of a given patent. By leveraging a multi-agent, tool-enhanced framework, PatentFinder decomposes the inherently complex task of patent assessment into specialized, manageable subtasks, each handled by tailored agents with specialized tools. As illustrated in Figure1, PatentFinder consists of five core LLM agents: Planner – assigns subtasks to specialized agents and compiles their outputs into a comprehensive infringement report, Sketch Extractor – identifies key molecular structures and converts them into Markush expressions, Substituents Matcher – verifies the substituent groups within Markush expressions relative to the target molecule, Requirements Examinator – evaluates if the query molecule meets the patent’s structural requirements, and Fact Checker – ensures accuracy by cross-verifying prior agents’ results against original claims and performing rollbacks if necessary. These agents independently engage with neural network models and heuristic tools to tackle specific subtasks, ultimately synthesizing a comprehensive patent infringement report.
The major innovations of PatentFinder include: (1) Multi-Agent Framework for Interpretability: To provide a clear reasoning path for LLMs, the multi-agent architecture enables PatentFinder to produce intermediate outputs, which are aggregated into a detailed infringement report. This design enhances the transparency and interpretability of the reasoning process; (2) Neural Network Models for Molecular Analysis: We developed two specialized models – MarkushMatcher and MarkushParser – to help LLMs accurately extract substituent groups and recognize Markush structures; (3) MolPatent-240 Benchmark: To facilitate systematic evaluation, we hire domain experts to curate MolPatent-240, a molecular patent infringement benchmark comprising 240 molecule-patent pairs with diverse and complex Markush structures. This dataset reflects real-world patent complexities, constructing a robust foundation for assessing the efficacy of patent infringement algorithms.
@@
Finally, extensive experiments in this paper demonstrate that PatentFinder has achieved significant breakthroughs in the accuracy and interpretability of patent infringement assessments: (1) PatentFinder achieves state-of-the-art performance, significantly outperforming baseline approaches that rely solely on LLMs. Specifically, PatentFinder surpasses baselines by 13.8% in F1-score and 12% in accuracy, demonstrating its effectiveness in handling the nuanced demands of patent infringement assessment; (2) As illustrated in Figure2e, PatentFinder can autonomously generate comprehensive patent infringement reports. These reports integrate detailed molecular structure analyses and step-by-step reasoning, resulting in highly informative and easily interpretable outputs. The improvements in accuracy and interpretability represent a significant step forward in enhancing the usability of automated patent analysis solutions. Moreover, our approach of decomposing complex tasks into modular components, each managed by a specialized tool agent, holds potential for application in other workflows of scientific discovery.
@@ 2 Results and Discussion
PatentFinder introduces two critical designs to facilitate the assessment of small-molecule patent infringement: (1) a multi-agent framework based on LLMs, where various agents equipped with tailored tools address manageable subtasks, and (2) two neural network tool models fine-tuned for specific tasks within the patent analysis process, addressing the performance deficiencies of LLMs in these tasks. We conduct extensive experiments to demonstrate the effectiveness of these designs. For the assessment of PatentFinder’s multi-agent framework, we compare PatentFinder with state-of-the-art LLMs on small-molecule patent protection discrimination. For the fine-tuned tool models, we evaluate them on the tasks of Markush molecular structure matching and Markush image recognition. The experimental results demonstrate that these tool models effectively compensate for the deficiencies of general-purpose LLMs in Markush understanding tasks.
@@ 2.1.1 Experiment Settings
To demonstrate the effectiveness of PatentFinder for patent infringement assessment, we construct a patent protection benchmark – MolPatent-240. To illustrate the rationale behind our agent-based design, we compare PatentFinder with the direct prompting of leading LLMs, which are tasked with reasoning and determining whether a given molecule infringes on a patent without any additional guidance. The baselines include state-of-the-art large language models Gemini-1.5-Pro [37] (abbreviated as Gemini), Claude-3.5-Sonnet 111https://www-cdn.anthropic.com/de8ba9b01c9ab7cbabf5c33b80b7bbc618857627/Model_Card_Claude_3.pdf. (abbreviated as Claude), GPT-4o [34], and OpenAI-o1 222https://openai.com/index/introducing-openai-o1-preview. We leverage the Chain-of-Thought prompts [40] for all the baseline models to enhance their reasoning ability. As shown in Table 1, we explore three distinct schemes to assess these models for patent infringement assessment: (1) Patent PDF: In this scheme, the patent PDF file is directly provided to the LLM. This approach requires the model to have a built-in PDF parsing capability. (2) Patent Text + Markush Image: The model is fed both the textual content of the patent and the core Markush structure images claimed in the patent. This scheme requires the model to handle both textual and visual information. (3) Patent Text + Markush String: In this case, all Markush images in the patent are first parsed into extended SMILES strings using our MarkushParser model. These Markush strings, along with the patent’s textual content, are then provided to LLMs in text format, allowing us to utilize text-only models in this scheme. Since the Patent PDF and Patent Text + Markush Image schemes require LLMs to handle PDF parsing and image processing, only Gemini, Claude, and GPT-4o are used for these experiments.
@@ 2.1.2 Results & Observations
Experimental results in Table 1 demonstrate that PatentFinder significantly outperforms the direct patent infringement classification with LLMs, regardless of the input scheme adopted. Specifically, our multi-agent collaboration framework achieves a 12.5% improvement in F1-score and a 12% increase in accuracy compared to the best-performing baseline model (Gemini with patent text + Markush string).
@@
Interestingly, we find that some popular LLMs (e.g., GPT-4o and Claude-3.5-Sonnet) tend to adopt an overly conservative classification approach, frequently assuming that the query molecule is protected by the given patent. They exhibit a TPR that is significantly higher than their TNR, often by 50% or more. Notably, GPT-4o (patent text + Markush string) displays a TPR that exceeds its TNR by 69.3%. Given that MolPatent-240 has a roughly balanced positive-to-negative example ratio (1:1.3), this trend suggests that many leading LLMs tend to classify samples as positive indiscriminately, with some models performing worse than random guessing when distinguishing negative samples (TNR). In contrast, PatentFinder demonstrates robust performance in distinguishing both positive and negative samples. It achieves a much lower class bias (TPR-TNR=0.9%), showcasing balanced and accurate classification across the MolPatent-240 dataset.
@@ 2.1.3 Case Studies
In real-world molecular patent infringement assessment, an effective and trustworthy collaboration between models and human experts requires the model to produce human-readable infringement analysis reports. These reports should not only deliver a definitive judgment on whether infringement has occurred but also provide a comprehensive and detailed explanation of intermediate reasoning steps. To this end, we evaluate not only the accuracy of the infringement assessment but also the validity of reasoning steps across different methods. Specifically, we compare the intermediate outputs of baseline models and PatentFinder on a real example from the MolPatent-240 dataset. The results are illustrated in Figure2. For baseline models, we summarize the step-by-step reasoning and final decisions of Gemini, GPT-4o, and Claude, while for PatentFinder, we present the intermediate outputs from each agent and the final infringement report generated by the Planner based on these analyses.
As shown in Figure2b, directly using LLMs for patent protection determination often leads to hallucinations in intermediate reasoning steps, resulting in incorrect conclusions based on non-existent evidence. For instance, Gemini provides no intermediate analysis and arrives at an incorrect judgment after merely comparing the molecular skeletons to the Markush structure. GPT-4o attempts to match the skeleton against the Markush structure but misidentifies the substituent D1 and introduces a fictitious substituent B. Although Claude reaches the correct conclusion, its analysis erroneously identifies R[21] as a chlorine atom and fails to determine the values of three other substituents.
From the baseline results, two key observations emerge: (1) Only GPT-4o attempts a fine-grained comparison of all substituents, while the other two models deliver conclusions without detailed analysis. (2) LLMs exhibit limited capabilities in comparing chemical structures, with reasoning errors propagating through subsequent analysis steps.
@@ 2.2 Evaluation of MarkushMatcher
To showcase the ability of MarkushMatcher to deliver precise structural matching results, we curate a dataset comprising 1,000 samples for substituent extraction. The comparison focuses on four key metrics: (1) Accuracy: The proportion of predictions that perfectly match the true substituent results. Perfect matches are defined as having a Tanimoto [42, 43] similarity of 100%. (2) Validity: The proportion of valid predictions. A prediction is considered valid if it includes all substituents mentioned in the Markush structure and each predicted substituent corresponds to a legitimate molecular structure. (3) Average Tanimoto Similarity: The average Tanimoto Similarity between each predicted substituent and its corresponding ground truth substituent, computed using Morgan Fingerprints [44] for all substituents in a sample, the higher the better. (4) Levenshtein Distance: The distance between the predicted results and the ground truth is calculated directly at the string level, the lower the better.
@@
We compare MarkushMatcher with structure-matching baselines implemented using the rule-based – RDKit [45] and LLM-based – GPT-4o methods, respectively. As shown in Figure 3a-c, MarkushMatcher outperforms both baseline methods on the substituent groups prediction task. As a deterministic algorithm, the rule-based approach – RDKit struggles with complex substituent matching tasks. While effective for simple tasks involving single-link substituents, it performs poorly in more intricate scenarios. This limitation result in near-zero accuracy and validity scores on our test set (cf. Figure 3a,b), with unsatisfactory results for the other two metrics as well. Using LLMs alone does not yield satisfactory results either. Although GPT-4o is able to provide chemically valid solutions in many scenarios, which results in the high validity in Figure 3c, it fails to predict the correct substituent group values in most cases, leading to low similarity, high Leveshtein distance, and limited accuracy. Common errors include: a) Confusing the values of different substituents; b) Incorrectly identifying substituent positions on the Markush structure; and c) Misclassifying scaffold atoms as part of the substituents. MarkushMatcher outperforms the other two methods by a large margin, with an accuracy of 66.8%. In comparison with GPT-4o, MarkushMatcher increases the mean average Tanimoto similarity from 21.3% to 92.9% and has an 8.8 times lower average Levenshtein distance. The training details of MarkushMatcher are presented in Figure 3d, e. During the model’s training, the training loss converges eventually, with a steady increase in both accuracy and validity. In the training data of MarkushMatcher, monatomic substituents such as C, N, and O are the most frequently occurring, while the heavy atom counts of other groups follow an approximately bell-shaped distribution.
@@
We also find that in certain cases, even with the given Markush and molecular structures, it is impossible to determine a unique result without additional contextual information. For instance, the presence of exchangeable groups (cf. Figure3f) and adjacent groups (cf. Figure3g) in the Markush structures can introduce complexities. Specifically, if two substituent groups are connected to the same atom, the correlation between the group identifiers and the substructures may become ambiguous. Similarly, in cases where multiple substituent groups are adjacent, there could be various valid allocation schemes when specifying the group values.